OpenStudy (anonymous):

jack mows the yard in 7 hours. marilyn mows it in 6 hours. how long would it take if they mowed the lawn together?

OpenStudy (saifoo.khan):

\[\frac 17 + \frac 16 + \frac 1x\]Solve for x.

OpenStudy (anonymous):

i hate story problems! they always give me problems!

OpenStudy (anonymous):

give me a second and i'll see what answer i get!

OpenStudy (saifoo.khan):

\[x=\frac{42}{13}\]

OpenStudy (anonymous):

thats what i came up with too! just couldn't figure out how to write the equation! always my weakness with story problems!

OpenStudy (unklerhaukus):

are they sharing the lawn mower or do they have two?

OpenStudy (saifoo.khan):

lol^

OpenStudy (anonymous):

42/13 hrs